# CATArena: Evaluation of LLM Agents Through Iterative Tournament Competitions

### v1

@@ Abstract
Large Language Model (LLM) agents have evolved from basic text generation to autonomously completing complex tasks through interaction with external tools. However, current benchmarks mainly assess end-to-end performance in fixed scenarios, restricting evaluation to specific skills and suffering from score saturation and growing dependence on expert annotation as agent capabilities improve. In this work, we emphasize the importance of learning ability, including both self-improvement and peer-learning, as a core driver for agent evolution toward human-level intelligence. We propose an iterative, competitive peer-learning framework, which allows agents to refine and optimize their strategies through repeated interactions and feedback, thereby systematically evaluating their learning capabilities. To address the score saturation issue in current benchmarks, we introduce CATArena, a tournament-style evaluation platform featuring four diverse board and card games with open-ended scoring. By providing tasks without explicit upper score limits, CATArena enables continuous and dynamic evaluation of rapidly advancing agent capabilities. Experimental results and analyses involving both minimal and commercial code agents demonstrate that CATArena provides reliable, stable, and scalable benchmarking for core agent abilities, particularly learning ability and strategy coding.
[…]
@@ 1 Introduction
With the rapid evolution of agents powered by large language models (LLMs), their capabilities have far surpassed simple text generation. By actively invoking external tools, LLM agents have significantly expanded the boundaries of artificial general intelligence (AGI). These agents are now able to autonomously complete complex, multi-step tasks that are previously considered beyond their reach, such as developing software project (Manish, 2024; Hu et al., 2025b), intelligently performing strategic planning (Belle et al., 2025), and learning user preference (Gao et al., 2024).
Existing benchmarks mainly focus on end-to-end performance in specific tasks, such as code generation (Yang et al., 2024), AI research (Nathani et al., 2025), and GUI automation (Wang et al., 2024). These benchmarks provide detailed observations and analyses of LLM agents’ abilities within particular scenarios and have driven significant progress in the field. However, there are important limitations to these approaches. First, the scores obtained in these end-to-end benchmarks only reflect performance on specific tasks, whereas an agent’s overall capability is composed of multiple fundamental skills working together. Second, the absolute scores in these benchmarks, which are typically based on objective correctness, have an upper bound. As agents become increasingly powerful, maintaining and updating these benchmarks requires additional expert-level annotation, and the level of required expertise continues to rise. In light of these challenges, there is an urgent need for a quantifiable and continuously evolving benchmark that systematically measures and analyzes the fundamental sub-abilities of agents.
Previous research has shown that self-learning is an essential ability for agents to achieve human-level intelligence (Gao et al., 2025; Zhu et al., 2025). Beyond self-learning, agents, similar to humans, also engage in peer learning, which enables collective evolution through interactions and shared experiences (Liu et al., 2024). During this evolutionary process, agents receive feedback from their environment and continually improve themselves. This capacity for learning and adaptation is indispensable for LLM agents, as it prepares them for ongoing evolution and more complex challenges. To systematically evaluate this crucial ability, we propose an iterative peer-learning-based competitive framework for LLM agents. In each iteration, agents are required to revise and update their strategies based on the outcomes and policies observed in previous rounds of competition. After every update, the agent policy codes are executed and competed against each other, generating dynamic performance rankings. Through this peer-learning architecture, we gain valuable insights into the learning abilities of LLM agents.
Building on this peer-learning framework, we introduce CATArena (Code Agent Tournament Arena) [^1] Code of CATArena is available in https://github.com/AGI-Eval-Official/CATArena., which utilizes four open-ended, rankable games. These games, including both board games and card games, provide LLM agents with a peer-learning environment and unlimited upper bound for improvement. They enable agents to continually improve and compete, ensuring that the evaluation framework remains challenging as agent capabilities grow. Furthermore, our competitive arena is inherently extensible and can be readily adapted to other types of open-ended, rankable tasks, facilitating the assessment of core agent abilities in new domains. As agent capabilities continue to advance, CATArena can evolve by incorporating tasks with greater complexity and discrimination, thereby supporting ongoing evaluation without the need for expert-level human annotation.
In our experiments, we conduct comparative performance evaluations and data analysis conducted on our self-developed minimal code agent and state-of-the-art commercial code agents. CATArena consistently provides stable and reliable benchmarks for assessing both agent capabilities and the agentic potential of the underlying LLMs. Within the peer-learning framework, we design general scoring metrics to systematically assess the fundamental abilities of participating agents, including their learning ability. Our experiments demonstrate that the strategy coding tasks applied in CATArena are fundamentally different from traditional LLM reasoning tasks. This represents a novel evaluation dimension that has not been addressed in previous work. Additionally, we analyze characteristics of CATArena, demonstrating its reliability and extensibility as a benchmarking platform.
@@
  • Iterative Peer-learning-based Competitive Framework: We propose a novel framework that leverages iterative peer-learning and competition to evaluate the learning abilities of LLM agents. Agents continuously revise their strategies based on feedback and outcomes from previous rounds, aligning agent evolution with human evolution.
  • CATArena Benchmark: We introduce CATArena, a tournament-style benchmark for evaluating the basic capabilities of LLM agents using a diverse set of open-ended games, including board and card games. CATArena provides an unlimited upper bound for agent improvement and supports extensible evaluation across diverse, open-ended tasks.
  • Comprehensive Agent Evaluation: We design general and systematic evaluation matrices and conduct comparative experiments and analyses between our minimal code agent and state-of-the-art commercial agents, demonstrating the reliability, stability, and extensibility of the CATArena.
@@ 2 Related Work
Learning Ability.
Learning ability is crucial for LLM agents, as it enables continual adaptation and improvement in dynamic and complex environments. Recent studies have shown that self-learning methods, such as self-refinement (Madaan et al., 2023; Shinn et al., 2023), allow models to enhance their outputs through iterative feedback, while environmental feedback further supports continual learning (You et al., 2024). In addition to self-learning, peer-learning has also been increasingly recognized, with approaches encouraging agents to learn from others’ reasoning processes and shared experiences (Liang et al., 2024; Luo et al., 2025). These diverse learning mechanisms have led to notable advances in tasks such as code generation, complex reasoning, and collaborative problem-solving. In the context of LLM-driven agents, learning ability represents a critical capability that supports effective adaptation and enables agents to tackle increasingly complex tasks and evolving challenges (Zhu et al., 2025; Gao et al., 2025). Despite the progress, systematic evaluation of how agents learn from each other remains underexplored, highlighting the need for benchmarks that capture both self-learning and peer-learning abilities.
Evaluation on Agents.
Recent benchmarks primarily assess LLM-driven agents on end-to-end, task-specific abilities. Code-based evaluations such as GitTaskBench (Ni et al., 2025), SUPER (Bogin et al., 2024), ProjectEval (Liu et al., 2025), SWE-PolyBench (Rashid et al., 2025), RedCode (Guo et al., 2024), SWT-Bench (Mündler et al., 2025), InfiAgent-DABench (Hu et al., 2024), and DA-Code (Huang et al., 2024) focus on large-scale software development, code security, bug fixing, and data science tasks. Other works extend agent evaluation to research (Du et al., 2025), real-world tool use (Yao et al., 2024), and assistant scenarios (Mialon et al., 2024). While some benchmarks explore agent-vs-agent evaluation (Zhuge et al., 2024), most rely heavily on human annotation and objective correctness, leading to upper bounds and saturation as agent capabilities advance.
Open-ended Tasks.
To address these limitations, recent work has leveraged open-ended, rankable tasks. Benchmarks such as GameBench (Costarelli et al., 2024), lmgame-Bench (Hu et al., 2025a), GAMEBot (Lin et al., 2024), and card game evaluations (Wang et al., 2025) assess LLMs’ strategic reasoning through diverse games. Frameworks like Game Reasoning Arena (Cipolina-Kun et al., 2025), GVGAI-LLM (Li et al., 2025), ZeroSumEval (Alyahya et al., 2025), TextArena (Guertler et al., 2025), and MCU (Zheng et al., 2025) further extend evaluation to multi-turn reasoning, spatial adaptability, natural language interaction, and open-ended tasks. However, these benchmarks mainly focus on reasoning skills and do not systematically evaluate agents’ learning abilities or coding strategies. Our analysis shows that measuring learning ability and coding strategy is fundamentally distinct, and both are essential for advancing agent intelligence. Additionally, in human board game competitions, variant rules such as Chess960 (FIDE, 2023) and Six-plus Hold’em (sixplusholdem, 2025) are often introduced to reduce memorization and encourage creativity. Notably, these variant rules have received relatively less attention in model evaluation.
3 CATArena
Figure 1: Overview of the evaluation framework and CATArena. The evaluation framework adopts an iterative peer-learning based competitive process. In round 1, LLM agents develop initial strategies via coding. These strategies are matched in a tournament arena, producing rankings and logs. In each subsequent round, agents analyze previous codes and logs, refine their strategies, and compete again. CATArena includes four open-ended and rankable games to cover diverse settings. Based on the tournament results from all rounds, a carefully designed scoring matrix and evaluation scheme are used to robustly quantify various agent abilities.
3.1 Iterative Peer-Learning based Competitive Framework
As shown in Figure 1, we propose an iterative peer-learning-based competitive framework, where CATArena evaluates code agents through a two-phase workflow: initial strategy development and iterative improvement. The initial phase assesses each agent’s ability to independently implement a baseline strategy based on the game code, while the iterative phase focuses on the agent’s learning ability.
Initial Development (Round 1). In this stage, each agent receives the game code and a sample AI implementation. Without external guidance, each agent must develop its own strategy to participate in the tournament. This phase primarily examines the agent’s strategy coding ability and establishes a baseline for subsequent evaluation.
Iterative Improvement (Rounds $n>1$). After the first round, all strategies submitted are evaluated through a tournament (round-robin format for symmetric games, batch-based competition for asymmetric games). Comprehensive competition logs are generated, recording rankings, win counts, and move histories for all matches. In subsequent rounds, agents are provided with the game code, previous round submissions from all participants, and these detailed logs. Agents must analyze these resources from previous rounds to adapt and improve their own strategies. This phase assesses the agent’s learning ability through repeated cycles of analysis and refinement.
This iterative evaluation framework of CATArena enables a granular assessment of both basic coding skills and advanced learning capabilities, supporting a robust and scalable measurement of code agent performance.
Table 1: Overview of game arenas and representative variants in CATArena.
Game Symmetry Type Players Variant
Gomoku Board 22 Forbidden points; dual three-in-a-row
Texas Hold’em Card $\geq 8$ Card removal; swapped hand ranks
Chess Board 22 Chess960; forbidden/special moves
Bridge Card 44 Card exchange
    For Bridge, symmetry is defined by assigning identical agent strategies to both teammates.
3.2 Games and Variants
Building on the tournament-based evaluation framework, CATArena deploys four distinct game arenas, each selected to test the strategic reasoning and coding capabilities of code agents across varying levels of complexity and interaction patterns. These arenas include competitive and cooperative settings, as well as symmetric and asymmetric game structures, thus enabling a diverse analysis of agents’ strategy coding abilities and learning patterns.
In addition to standard rules, each game is extended with thoughtfully designed variants that introduce novel or altered mechanics, inspired by real-world adaptations such as Fischer Random Chess (Chess960) (FIDE, 2023). Like human competition, the variant rules encourage strategy generalization and penalize rote memorization, as most models are trained on card and board game data. Table 1 provides an overview of the selected games and their respective variants.
3.3 Tournament Format and Scoring System
After the completion of all $N$ development rounds, CATArena conducts a comprehensive tournament to quantitatively evaluate agent strategies and compute performance metrics. A total of $T$ agent models participate, each contributing strategies in every round of development. Tournament formats are tailored to game types: for symmetric games, all strategies engage in a round-robin cycle, ensuring exhaustive pairwise competition; for asymmetric games such as Texas Hold’em, strategies are grouped into batches and compete in multi-agent matches. To mitigate randomness, all matches are repeated multiple times, and results are averaged for robust evaluation.
Scores are recorded in a scoring matrix $W\in\mathbb{R}^{(TN)\times(TN)}$, where $W_{i,j}^{n,m}\in[0,1]$ denotes the score obtained by agent $i$’s strategy in round $n$ against agent $j$’s strategy in round $m$. When $n=m$, the notation simplifies to $W_{i,j}^{n}$; similarly, when $i=j$, it is denoted as $W_{i}^{m,n}$. This scoring system enables fine-grained, quantitative analysis of agent performance in both individual and iterative development stages. For asymmetric games, pairwise results are not feasible; instead, batch-based tournaments are used and the score matrix records the win rates of multi-agent matches. The tournament format is provided in the Appendix A.
3.4 Evaluation Metrics
Based on the scoring matrix $W$, we design a set of evaluation metrics to quantitatively assess the key capabilities of code agents. Specifically, our metrics are constructed to measure three core capabilities: strategy coding, learning, and generalizability. In the following sections, we define these metrics using symmetric games as examples. For asymmetric games, the evaluation principles remain consistent. The calculation of the scoring matrix $W$ is provided in Appendix B.
Strategy Coding.
Strategy coding measures the agent’s fundamental ability to abstract game strategies into reproducible algorithms and implement them as executable code, which is fundamentally different from general reasoning and strategic planning abilities. In CATArena, this metric evaluates how effectively an agent can independently develop a baseline strategy for the game environment and compete against other agents in the initial development stage.
For each agent $i$, strategy coding is quantified by the average score obtained against all other agents in the first round:
$S_{i}=\mathrm{avg}_{j\neq i}(W_{i,j}^{1}).$
This metric serves as the foundational benchmark for code agent evaluation in CATArena.
Learning Ability.
The learning capability of a code agent captures its ability to leverage historical information and opponent behaviors to improve its own performance.
Global Learning assesses an agent’s overall improvement in strategy quality. This metric evaluates the relative performance of agent $i$’s strategies against all strategies from all agents and rounds, and measures the average progress made compared to its initial baseline. It serves as the primary indicator of learning ability.
Formally, for agent $i$, global learning is defined as:
$L_{i}=\mathrm{average}_{n=2}^{N}\left(G_{i}^{n}-G_{i}^{1}\right),$
where $G_{i}^{n}$ represents the global performance of agent $i$’s strategy from round $n$:
$G_{i}^{n}=\mathrm{average}_{(i,n)\not=(j,m)}\left(W_{i,j}^{n,m}\right).$
This metric captures the agent’s ability to learn and adapt over multiple rounds, reflecting its progress in a comprehensive competitive landscape.
Counter-Adaptation measures an agent’s targeted learning ability, reflecting its capacity to achieve improved results against opponents in successive rounds. For agent $i$, the counter-adaptation score is defined as the average improvement in scores against other agents from round n1n-1 to round $n$ $(n\geq 2)$:
$C_{i}=\mathrm{average}_{n=2}^{N}\left(A_{i}^{n}-B_{i}^{n-1}\right),$
where the advance score $A_{i}^{n}$ and base score $B_{i}^{n-1}$ are defined as:
$A_{i}^{n}=\mathrm{average}_{j\neq i}\left(W_{i,j}^{n,n-1}\right),B_{i}^{n-1}=\mathrm{average}_{j\neq i}\left(W_{i,j}^{n-1}\right).$
Here, $A_{i}^{n}$ represents agent $i$’s average performance in round $n$ against the strategies submitted by other agents in the previous in round (n1n-1). The base score $B_{i}^{n-1}$ denotes agent $i$’s average performance in round n1n-1 against those same opponents. This comparison isolates the agent’s targeted adaptation from one round to the next.
Self-improvement evaluates an agent’s capacity to genuinely enhance its strategies over successive rounds of development. This metric reflects whether newly developed strategies can consistently outperform the agent’s own previous versions.
We quantify self-improvement by calculating the Pearson correlation Pearson (1896) between the round index and the agent’s average scores across rounds. For agent $i$, the self-improvement score is defined as
$\mathrm{SI}_{i}=\mathrm{Pearson}\left([1,\cdots,N],\ [S_{i}^{1},\cdots,S_{i}^{N}]\right).$
Here, $S_{i}^{n}$ denotes the average score of agent $i$’s strategy from round $n$ against its own strategies from other rounds
$S_{i}^{n}=\mathrm{average}_{m\neq n}\left(W_{i}^{n,m}\right).$
A higher self-improvement score indicates a stronger ability to iteratively refine and upgrade strategies throughout the development process.
Generalizability.
Generalizability measures an agent’s ability to comprehend and adapt to novel or altered game rules that differ from those encountered during training or prior experience. This metric specifically evaluates the agent’s capacity to generalize beyond previously seen environments, focusing on handling new or modified scenarios. For agent $i$, the generalizability score is defined as:
$U_{i}=B_{i}^{1;\mathrm{Variants}}-B_{i}^{1;\mathrm{Standard}},$
where $B_{i}^{1;\mathrm{Variants}}$ and $B_{i}^{1;\mathrm{Standard}}$ denote the base scores of agent $i$ in the first round under variant and standard rule settings, respectively. A higher value of $U_{i}$ indicates stronger generalizability, reflecting the agent’s ability to effectively develop and apply strategies for previously unseen tasks.

### v2

# CATArena: Evaluating Evolutionary Capabilities of Code Agents via Iterative Tournaments

@@ Abstract
Current evaluation for Large Language Model (LLM) code agents predominantly focus on generating functional code in single-turn scenarios, which fails to evaluate the agent’s capability for continuous code optimization and multi-turn iterative development. To bridge this gap, we introduce CATArena, a framework designed to evaluate the evolutionary capabilities of code agents via iterative tournaments. Agents engage in multi-turn tournaments and continuously refine their code through self-reflection and peer-learning based on comprehensive execution feedback. For evaluation, we propose a dual-metric system to decouple static generation proficiency from evolutionary potential. Extensive experiments reveal that an agent’s evolutionary potential is not strictly correlated with its initial proficiency. Our analysis further reveals that current agents struggle to concurrently leverage both peer-learning and self-reflection for effective performance gains. Furthermore, the results validate CATArena’s high extensibility and resistance to variance tasks, establishing it as a continuous and reliable standard for assessing the evolutionary capability of LLM code agents.  [^1] The official website of CATArena is https://catarena.ai.
Code Agents, Evaluation, Evolutionary Capabilities
@@ 1 Introduction
The rapid evolution of Large Language Models (LLMs) has catalyzed the emergence of code agents. Unlike passive chat assistants, code agents actively interact with dynamic digital workspaces, manipulating file systems and utilizing toolchains to orchestrate complex software development workflows (Dong et al., 2025; Dearing et al., 2025; Xu et al., 2025c). To assess these expanding capabilities, the evaluation landscape has evolved from simple method-level generation (Chen et al., 2021) to repository-level engineering scenarios (Xu et al., 2025a, b). Performance is predominantly quantified by static metrics such as task success rate (Yang et al., 2024) within a single-turn development process. While some emerging benchmarks incorporate non-functional metrics such as security or code complexity, these remain static checks rather than dynamic objectives.
These static and correctness-centric benchmarks impose two critical limitations on advancing agent intelligence. First and foremost, existing protocols suffer from the absence of iterative evolution. Real-world software development is inherently incremental and collaborative (Bennett, 1996): human engineers continuously refine their solutions by analyzing execution logs and learning from peer reviews (Liu et al., 2024). Crucially, this evolutionary mechanism is equally indispensable for autonomous agents. However, by restricting evaluation to a single-turn mode, current frameworks isolate agents from the dynamic feedback loops essential for optimization and evolution. Agents are prevented from rectifying runtime errors through self-debugging or optimizing strategies by observing superior peer implementations (Gao et al., 2025; Zhu et al., 2025a). Consequently, current benchmarks capture only a static snapshot of the agent’s baseline competency, failing to evaluate its dynamic capacity for self-reflection and peer-learning over time.
Furthermore, this lack of iterative depth leads to an underemphasis on code quality optimization. Since evaluation is confined to a single turn, current frameworks predominantly rely on binary functional correctness metrics like pass@k (Chen et al., 2021; Austin et al., 2021), failing to incentivize agents to refine code for computational efficiency or robustness. In large-scale software engineering, a functionally correct but inefficient algorithm can lead to unacceptable system latency or energy consumption (Dearing et al., 2025). However, standard benchmarks treat a brute-force solution as equivalent to an optimized expert implementation. This equivalence masks the performance gap between agents that merely solve a problem and those that engineer superior solutions, causing metrics to saturate without capturing the potential for unbounded evaluation.
To bridge this gap, we introduce CATArena (Code Agent Tournament Arena), a dynamic evaluation framework designed to assess the evolutionary capabilities of code agents through iterative refinement. Unlike traditional static benchmarks, CATArena immerses agents in a multi-turn development loop, mimicking the continuous improvement process of human engineers. Within this interactive workspace, agents must go beyond generating functional code. They are required to analyze execution feedback and, crucially, leverage peer solutions and opponent strategies to iteratively optimize their performance. To ensure an unbounded evaluation, we instantiate six diverse tasks with high performance ceilings and design scoring mechanisms that reward continuous optimization. These tasks are categorized into two types based on the feedback signal: two objective optimization tasks, where agents optimize against fixed metrics (e.g., time complexity, throughput); and four competitive tasks, a novel introduction where agents adaptively optimize their strategies against dynamic opponents (e.g., in strategy games, network attack and defense). This dual-track design allows the framework to seamlessly integrate diverse task types while effectively measuring an agent’s potential to evolve from basic correctness to expert-level excellence.
We conduct extensive experiments using a multi-round iterative process, evaluating both self-developed minimal agents powered by various LLMs and state-of-the-art (SoTA) commercial code agents. Our analysis reveals that evolutionary capability is a distinct dimension of intelligence, independent of single-turn generation capability. We observe that agents with high initial pass rates do not necessarily possess strong evolving abilities. Conversely, some agents demonstrate low-start, high-growth trajectories, effectively leveraging feedback to surpass initially superior competitors. Furthermore, by analyzing the optimization trajectories, we identify distinct evolutionary patterns in how agents utilize peer solutions and execution logs. Ultimately, these results demonstrate that the iterative framework enables agents to overcome the performance plateaus typical of single-round development, validating the necessity of unbounded evaluation for the next generation of intelligent agents.
@@
  • Evolutionary Capability Evaluation: To the best of our knowledge, we are the first to identify and formalize the evolutionary capability of code agents, aligning with the iterative nature of real-world engineering.
  • CATArena Framework: We introduce CATArena, a tournament-based evaluation framework that models the development process as a multi-turn evolutionary loop. Driven by iterative feedback, it quantifies code refinement via self-reflection and peer-learning, transcending static, single-turn evaluation.
  • Comprehensive Agent Analysis: We demonstrate that evolutionary capability is orthogonal to static generation performance and analyze its temporal dynamics across tasks. Our extensible experiments further highlights CATArena’s extensibility with new environments, revealing that SoTA agents still faces significant challenges on complex tasks.
@@ 2 Related Work
Figure 1: Overview of the CATArena Framework. (1) Initial Development: Agents receive task specifications and generate initial solution code $C^{(1)}$. (2) Execution Feedback: Submissions are evaluated in either competitive or objective tracks. Results are aggregated into a structured set containing execution logs, leaderboards, and peer solutions. (3) Iterative Evolution: Agents engage in self-reflection and peer-learning to refine their code $C^{(n)}$ for the subsequent rounds.
Evolutionary Mechanisms. Evolutionary capability is pivotal for agents to adapt in dynamic environments. For LLM-based agents, this manifests as in-context evolution, where agents optimize strategies using historical context and execution feedback without parameter updates (Zhu et al., 2025a). Recent research highlights two primary mechanisms driving this evolution: self-reflection and peer-learning. Self-reflection methods (Madaan et al., 2023; Shinn et al., 2023) enhance outputs through iterative self-correction based on environmental feedback (You et al., 2024). Peer-learning approaches encourage agents to distill insights from the reasoning processes and solutions of others (Liang et al., 2024; Luo et al., 2025). While these mechanisms have advanced code generation and reasoning, the systematic evaluation of an agent’s evolutionary capability remains underexplored, particularly regarding how agents leverage feedback to evolve over time in competitive coding scenarios. CATArena fills this void as the first to assess the evolutionary capability of code agents.
Evaluation of Code Agents. Recent benchmarks have scaled evaluation to simulate complex, real-world software engineering scenarios. In the domain of repository-level development, frameworks such as GitTaskBench (Ni et al., 2025), SWE-PolyBench (Rashid et al., 2025), and SWT-Bench (Mündler et al., 2025) challenge agents with large-scale maintenance and bug fixing, complementing established standards like SWE-bench (Yang et al., 2024). Beyond general engineering, specialized evaluations have emerged for diverse domains: RedCode (Guo et al., 2024) and SecBench (Lee et al., 2025) focus on code security; InfiAgent-DABench (Hu et al., 2024) and DA-Code (Huang et al., 2024) target data science workflows; while other works assess research automation (Du et al., 2025) and web development (Xu et al., 2025b). Despite this extensive coverage, the evaluation methodology remains predominantly static. Performance is typically quantified by binary success metrics such as Pass@k (Chen et al., 2021) and task success rate (Yang et al., 2024). Although some benchmarks incorporate non-functional metrics like code complexity (Siddeeq et al., 2025) or test coverage (Chen and Jiang, 2025), these are applied as static checks.
To overcome the limitations of static metrics, research has expanded into open-ended tasks. In the realm of strategic reasoning, benchmarks like GameBench (Costarelli et al., 2024) and TextArena (Guertler et al., 2025) assess agents through multi-turn gameplay. Similarly, for code optimization, datasets such as OIBench (Zhu et al., 2025b) and SWE-bench (Yang et al., 2024) provide open-ended challenges for algorithmic reasoning and engineering performance. However, current evaluations for these tasks are predominantly single-turn, treating them as static problems rather than opportunities for continuous optimization. This restriction prevents code agents from demonstrating their full potential through iterative refinement. CATArena bridges this gap by incorporating these open-ended tasks into an evolutionary framework, rigorously testing the limits of code agents.
3 CATArena Framework
In this section, we detail the CATArena framework, a universal and robust evaluation paradigm designed to shift the assessment of code agents from static correctness to dynamic evolutionary capability.
3.1 Iterative Evolution via Tournament Feedback
Unlike traditional benchmarks that treat coding as a static, single-turn task, CATArena abstracts the evaluation process into a continuous loop of development, execution, and evolution. As illustrated in Figure 1, this paradigm is structured into two primary phases bridged by an execution feedback mechanism.
Table 1: Overview of tasks in CATArena. The framework instantiates six diverse tasks. Crucially, the feedback mechanism provides agents with population-level data, including full tournament logs and peer solutions, rather than just individual execution results, fostering group evolution.
Category Task Domain Players Feedback Content
Competitive Gomoku Board Game 2 Adversarial Match Logs & Peer Solutions
Texas Hold’em Card Game $\geq 8$
Chess Board Game 2
Bridge Card Game 4
Objective OIBench (Zhu et al., 2025b) Algorithm Optimization 1 Deterministic Efficiency Metrics & Peer Solutions
SWE-Perf (He et al., 2025) Software Optimization 1
Initial Development (Round 1). The process begins with the initial development phase, which establishes the single-turn development capability of the code agent. In this stage, the agent receives a task specification $\mathcal{T}$, ranging from game rules to algorithmic problems, along with a minimal functional example. Operating without external guidance or historical data, the agent must comprehend the requirements and implement an initial solution $C^{(1)}$.
Execution Feedback. Upon receiving the code submission, CATArena initiates the execution feedback process to drive the transition from static code generation and dynamic evolution. The system executes the submitted code within a sandboxed evaluation arena to generate quantitative performance scores. This phase aggregates execution results into a structured feedback set containing comprehensive tournament execution logs, a global leaderboard, and peer solutions. While the specific metrics and feedback modalities vary depending on the task type (as detailed in Section 3.2), the fundamental function remains consistent in providing the necessary signals for evolution.
Iterative Evolution (Round $n>1$). In the iterative evolution phase, the goal of agents shifts from generation to optimization. Agents are provided with the task specification and the execution feedback derived from the previous round. By analyzing this information, agents are tasked to identify weaknesses in their previous implementations via self-reflection and assimilate superior approaches through peer-learning. The agent then synthesizes these insights to generate an evolved solution $C^{(n)}$. This cycle repeats for $N$ rounds, simulating the real-world engineering process where solutions are continuously refined based on testing and analysis of external implementations.
3.2 Task Categorization and Instantiation
To comprehensively evaluate the evolutionary potential of code agents, CATArena instantiates six diverse coding tasks. As demonstrated in Table 1, these tasks are categorized into two distinct types based on the nature of the feedback source, mirroring the duality of real-world software engineering where development is driven either by adversarial dynamics or distinct system efficiency targets: Competitive Tasks and Objective Optimization Tasks.
Competitive Tasks. This category models scenarios where performance is inherently relative. Unlike static benchmarks, the feedback mechanism here is dynamic: a solution’s effectiveness depends entirely on the strength of its opponents. As the population evolves, the difficulty of winning increases, creating a non-stationary evaluation target that compels agents to continuously adapt to new counter-strategies. To capture this, we employ a Round-Robin tournament structure where the code implementations submitted by agents are executed against each other in pairwise or multi-agent matches. Crucially, the feedback provided to agents is not limited to their own match results. It encompasses the full tournament logs and peer solutions from the entire population. We instantiate four specific arenas: Gomoku, Texas Hold’em, Chess, and Bridge. This diversity, covering both symmetric perfect-information games and asymmetric imperfect-information settings, rigorously tests the agent’s ability to analyze complex interaction dynamics and refine strategies against an adaptive field.
Objective Optimization Tasks. This category models scenarios where the primary goal is to optimize a static objective function. Here, the feedback is quasi-static. While the evaluation metrics are deterministic, the competitive standard is continuously raised by the population. Agents must treat functional correctness merely as a starting point, striving to approach theoretical limits to outperform peer solutions. We incorporate two specific benchmarks. First, OIBench (Zhu et al., 2025b) evaluates algorithmic efficiency using Time AUC, a metric that quantifies the volume of test cases passed within constrained time limits. Second, SWE-Perf (He et al., 2025) assesses repository-level engineering by challenging agents to optimize authentic GitHub codebases. It measures the performance speedup relative to the original implementation, highlighting the gap between functional code and expert-level optimization.
3.3 Formalization and Metrics
To provide a standard of unified evaluation across diverse tasks, we construct a global performance matrix $W$. Based on this matrix, we derive two core metrics to comprehensively assess both the agent’s static coding capability and its evolutionary potential.
Performance Matrix Construction. Let $N$ be the total number of rounds and $M$ be the number of participating agents. We define the performance matrix $W\in\mathbb{R}^{(M\times N)\times(M\times N)}$, where $W_{i,j}^{n,m}$ denotes the performance score of Agent $i$ (from Round $n$) evaluated against Agent $j$ (from Round $m$).
  • For competitive tasks, $W_{i,j}^{n,m}$ represents the win rate of agent $i$’s code in round $n$ against agent $j$’s code in round $m$. For symmetric settings, this is derived from pairwise round-robin matches. For asymmetric settings (e.g., Texas Hold’em), we utilize batch-based tournaments where scores reflect the win probability in the tournament.
  • For objective optimization tasks, since agents operate independently. Thus, we simplify the notation $W_{i,j}^{n,m}$ to $W_{i}^{n}$, representing the scalar performance score of agent $i$ at iteration $n$. Specifically, for OIBench, we adopt their Time AUC metric (Zhu et al., 2025b), which quantifies the area under the curve of passed test cases against varying time constraints, normalized by the maximum achievable area. For SWE-Perf, we utilize the performance ratio (He et al., 2025), defined as the computational speedup of the optimized code relative to the original baseline, normalized by a target expert reference.
Detailed formalization of the matrix $W$ in CATArena is provided in Appendix A. To mitigate randomness in game outcomes, every entry in the scoring matrix $W$ is estimated by repeated matches for fixed code submissions. Detailed repetition settings for each task are provided in Appendix B.
Basic Performance $S_{\text{base}}$. This metric evaluates the agent’s static coding ability in a traditional single-turn setting. It quantifies the relative performance of the agent’s initial submission $C^{(1)}$ within the Round 1 population:
$S_{\text{base},i}=\frac{1}{M-1}\sum_{j\neq i}W_{i,j}^{1,1}.$ (1)
For objective tasks, this simplifies to the score of the first round $W_{i}^{1}$. $S_{\text{base}}$ serves as a baseline to reflect the agent’s zero-shot task-solving proficiency before any feedback is introduced.
Evolutionary Capability $S_{\text{evo}}$.
This metric is the primary indicator of evolutionary capability. Unlike static metrics, it quantifies the rate of improvement driven by feedback.
We first calculate the global performance score $G_{i}^{n}$ for each round $n$. For competitive tasks, $G_{i}^{n}$ represents the relative strength of Agent $i$’s Round $n$ code when evaluated against the entire population of code versions ($M\times N$) produced across all rounds:
$G_{i}^{n}=\frac{1}{MN-1}\sum_{(j,m)\neq(i,n)}W_{i,j}^{n,m}.$ (2)
We use global average score for $G_{i}^{n}$ and give a brief explanation in Appendix C about the stability of this metric. For objective tasks, $G_{i}^{n}$ is simply the score $W_{i}^{n}$.
$S_{\text{evo},i}$ is derived as the slope $k_{i}$ of the linear fit $G_{i}^{n}\approx k_{i}\cdot n+b_{i}$ using the Ordinary Least Squares (OLS) method:
$S_{\text{evo},i}=\frac{\sum_{n=1}^{N}(n-\bar{n})(G_{i}^{n}-\bar{G}_{i})}{\sum_{n=1}^{N}(n-\bar{n})^{2}},$ (3)
where $\bar{n}$ and $\bar{G}_{i}$ are the average round number and average global performance, respectively. A positive $S_{\text{evo}}$ indicates effective adaptation, with higher values signifying a faster rate of evolution.
We employ OLS method primarily because development failures might occur in any round due to inherent limitations in their robustness. Compared to simpler alternatives such as total improvement (i.e., $G^{N}-G^{1}$), which are highly sensitive to single-round outliers, OLS provides a more robust estimation of the overall growth trend by mitigating the impact of transient failures. Importantly, we regard development failures as a failure of the evolutionary process itself, and these instances are explicitly factored into the OLS calculation to reflect the agent’s inability to maintain stable progress. Furthermore, as our experiments focus on the initial active evolution phase ($N\leq 4$), a linear approximation effectively captures the rate of evolution before performance saturation.